OpenStudy (clamin):

PLEASE HELP!! i keep figuring this out but im kind of confused 1. Calculate the number that you need to add to each side of the equation w^2 - 3w = 350 to create a perfect square trinomial. 2. Factor the trinomial. 3. Now that you've factored the equation, find the square root of each side and solve for w.

OpenStudy (clamin):

this is what i got for the first question

OpenStudy (kittiwitti1):

What I think the whole problem is basically saying: 1. complete the square 2. factor it out... 3. (so you can) solve for w

OpenStudy (clamin):

this one i got for the last question @kittiwitti1

OpenStudy (kittiwitti1):

...what Can you show me the entire process? lol

OpenStudy (clamin):

ok hold on

OpenStudy (johnweldon1993):

You were good until step 5 You brought back the w^2 out of nowhere...but yes, we'll see your whole process!

OpenStudy (kittiwitti1):

sorry, I was afk but I sent you a PM regarding the steps to take after finding the\[(\frac{b}{2})^{2}\]

OpenStudy (kittiwitti1):

Also, Step two is missing the +(-3/2)^2 on the right side of the equation. Or you could eliminate it entirely since Step three does that. http://prntscr.com/a75r56

OpenStudy (anonymous):

Remember that\[ \sqrt{\left(w-\frac 32\right)^2} = \left|w-\frac 32\right| \]

OpenStudy (kittiwitti1):

Oh, right. The answer will be both positive and negative\[\pm(\text{answer})\]

OpenStudy (kittiwitti1):

Thanks @wio

OpenStudy (anonymous):

So you have to use a \(\pm\) on the other side to get rid of it.

OpenStudy (clamin):

this one is the second process

OpenStudy (kittiwitti1):

ok so I did this:\[w^{2}-3w+(-\frac{3}{2})^{2}=350+(-\frac{3}{2})^{2}\rightarrow (x-\frac{3}{2})^{2}=350+\frac{9}{4}\]\[(x-\frac{3}{2})^{2}=\frac{1400}{4}+\frac{9}{4}\rightarrow x-\frac{3}{2}=\pm\sqrt{\frac{1409}{4}}\]\[x=\frac{3}{2}+\frac{\sqrt{1409}}{2}\rightarrow \frac{3\pm\sqrt{1409}}{2}\]

OpenStudy (kittiwitti1):

Is this roughly along the lines of how you solved it?

OpenStudy (kittiwitti1):

http://prntscr.com/a75vax

OpenStudy (clamin):

how did you get 1400?? did you mutiply it by 9??

OpenStudy (johnweldon1993):

\[\large 350 + \frac{9}{4}\] Find the common denominator \[\large \frac{1400}{4} + \frac{9}{4}\]

OpenStudy (kittiwitti1):

^ what he said, yes

OpenStudy (kittiwitti1):

\[350\times\frac{4}{4}=\frac{1400}{4}\]

OpenStudy (kittiwitti1):

Everything good so far?

OpenStudy (clamin):

yes

OpenStudy (kittiwitti1):

Okay. Can you do it from here or should I keep going?

OpenStudy (clamin):

yea i think i can dot it

OpenStudy (clamin):

does this makes sense the width is \[\frac{ 3\pm \sqrt{1409} }{ 2 }\] ??

OpenStudy (kittiwitti1):

I believe you have to do both plus and minus and get two answers. Or only one of them work. What was the original context of the question?

OpenStudy (clamin):

it asks: what are the maximum dimensions of the stage that malcolm can build? w=_________ l=__________

OpenStudy (kittiwitti1):

Hmm... I think you have to input two separate equations into your calculator:\[\frac{3+\sqrt{1409}}{2}\text{ and }\frac{3-\sqrt{1409}}{2}\]If you get a negative answer then that would be the one that doesn't work, since measurements like this need to be positive.

OpenStudy (clamin):

ok

OpenStudy (kittiwitti1):

Good luck ♣